OpenStudy (anonymous):

Find the equation of a line which passes through point (6, 8) and is perpendicular to the line formed by the equation y = 6x + 1. Express it in slope-intercept form.

OpenStudy (amistre64):

perp lines are 90 degrees from each other; and they have the inherent property that when you multiply their slopes together you get -1

OpenStudy (anonymous):

ok here he goes.

OpenStudy (anonymous):

haha

OpenStudy (amistre64):

one simple way to get a perp line is to swith out the numbers on your x and y variables; and negate a sign

OpenStudy (amistre64):

y = 6x + 1 6y = x + n; negate one of them like this 6y = -x + n ; we can solve for y by dividing out the 6 now y = (-x+n)/6

OpenStudy (amistre64):

the "n" is determined by the point being used; fill in the points and solve for n

OpenStudy (amistre64):

y = (-x + n)/6; (6,8) 8 = (-6 + n)/6 48 = -6 + n 42 = n y = (-x+42)/6 whould do it

OpenStudy (amistre64):

42 is wrong, I forgot how to add lol

OpenStudy (amistre64):

y = (-x+54)/6

OpenStudy (anonymous):

or you can try this. we know (as sensei said) that the slope if \[-\frac{1}{6}\]now use "point-slope" formula \[y-y_1=m(x-x_1)\] with \[m=-\frac{1}{6}, x_1=6, y_1=8\]

OpenStudy (amistre64):

another way to write it is: y = (-1/6)x + 54/6 = (-1/6)x + 9

OpenStudy (anonymous):

get \[y-8=-\frac{1}{6}(x-6)\] \[y-8=-\frac{1}{6}x+1\] \[y=-\frac{1}{6}x+9\]